OpenStudy (theviper):

Please Help :) Solve the Inequality:- \[\Huge{\color{orange}{\frac{6x-5}{4x+1}<0}}\] :D

mathslover (mathslover):

Ok so here i go : \[\Huge{\color{orange}{\frac{6x-5}{4x+1}<0}}\] Multiply both sides by 4x + 1 ,,, can we start like this @experimentX ?

OpenStudy (anonymous):

solve this \[\Huge{\color{orange}{\frac{6x-5}{4x+1} = 0}}\] \[\Huge \color{orange}{ a x = b }\] Put \[\Huge \color{orange}{ a x < b }\]

OpenStudy (experimentx):

yep!!

mathslover (mathslover):

ok thanks so wht r u getting @TheViper after multiplying both sides by 4x + 1

OpenStudy (theviper):

6x-5<0

mathslover (mathslover):

Ok correct : so add 5 both sides. . wht r u getting ?

OpenStudy (theviper):

Ok now I can solve thanx everyone:)

mathslover (mathslover):

nice to hear .. good work @TheViper

OpenStudy (theviper):

^_^

OpenStudy (anonymous):

is the answer -1/4 < x < 5/6

OpenStudy (theviper):

@mathslover @experimentX Brothers, attention plz , umm...., as we don't know the value of 'x' so we also don't know the value of the value of '4x+1' that it is negative or positive.So we can't multiply it with 0 becoz if it is negative it will reverse the sign. Am I correct??

OpenStudy (experimentx):

hmm ... can you plug that inequality into wolf?

OpenStudy (theviper):

Plz big bro u plug I don't know very much to plug any ques in wolfy :)

OpenStudy (theviper):

I think ans should be this:-\[\Large{x \epsilon(-\frac{1}{4},\frac{5}{6})}\]

OpenStudy (experimentx):

on that's a complete unexpected result http://www.wolframalpha.com/input/?i=%286x%E2%88%925%29%2F%284x%2B1%29%3C0

OpenStudy (theviper):

I can't understand wht wolfy says, plz u tell :) @matricked has given the same answer as wolfy:)

OpenStudy (experimentx):

well ... see below, it does  cross multiply it and get the values. well i didn't know that either ...